OpenStudy (anonymous):

How do you solve this problem? 8/3x^3-5/12x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

b solve you mean?

OpenStudy (anonymous):

most likel;y differentiate

OpenStudy (anonymous):

with the cubic power over 3

OpenStudy (amistre64):

\[\frac{8}{3}x^3-\frac{5}{12}x\] like this equation?

OpenStudy (amistre64):

....i meant expression :)

OpenStudy (anonymous):

Exactly like that equation .

OpenStudy (amistre64):

and by solve; factor it?

OpenStudy (anonymous):

Well, I am supposed to find the sum or difference and put it in simple form, I dint see exactly what to factor in the expression.

OpenStudy (anonymous):

I dont kow why a teacher wuld ask you to factor that lol

OpenStudy (anonymous):

its pointless, all you can do is take out an x

OpenStudy (amistre64):

since its asum of differeing powers of x; there is no way to add it

OpenStudy (amistre64):

\[\frac{1}{3}x(8x^2 -\frac{5}{4})\] perhaps?

OpenStudy (amistre64):

maybe even take out a 1/4

OpenStudy (amistre64):

\[\frac{1}{12}x(32x^2-5)\]

OpenStudy (anonymous):

So you want me to distribute 1/12x ?

OpenStudy (amistre64):

i dont want you to mess with it; but i see nothing else to do to it :)

OpenStudy (amistre64):

\[\frac{8}{3x^3}-\frac{5}{12x}\] perhaps this is the problem?

OpenStudy (anonymous):

That is silimar one of my choices except it looks like 32-5x^2 over 12x^3

OpenStudy (anonymous):

Yes, That was the problem.

OpenStudy (amistre64):

...... now it can be both :)

OpenStudy (amistre64):

i typed cant, im sure i did. my personalitites are conspiring against me lol

OpenStudy (anonymous):

ohh hah , i was confused.

OpenStudy (amistre64):

\[\frac{8}{3x^2}*\frac{4}{4}-\frac{5}{12x}*\frac{x}{x}\] \[\frac{32}{12x^3}-\frac{5x}{12x^3}\] \[\frac{32-5x}{12x^3}\]

OpenStudy (amistre64):

that little 3x^2 is a typo; it should be 3x^3

OpenStudy (anonymous):

So you multiplied 4 on both sides and then x why? o:

OpenStudy (amistre64):

you gotta read thru the typos; but you get the basic principals right?

OpenStudy (amistre64):

the x^2/x^2 =1; so we change the way the fraction looks without changing its value; and that way we can get like bottoms

OpenStudy (amistre64):

4/4 = 1 as well

OpenStudy (anonymous):

Okay I understand , Thanks a lot I was confused for quite a while.

OpenStudy (amistre64):

:) practice helps. And no typos too...